A parent pushes a baby stroller from home to daycare along a level road with a force of 40 N directed at an angle of 30° below the horizontal. If daycare is 0.83 km from home, how much work is done by the parent?
Answer #1

Work-done is given by:

W = F.d

W = F*d*cos A

F = Force applied by parent = 40 N

d = displacement of baby from home to daycare = 0.83 km = 830 m

A = Angle between applied force and displacement = 30 deg

So,

W = 40*830*cos 30 deg

W = 28752.04 J

W = 2.9*10^4 J

Please Upvote.

Comment below if you've any query.